OpenStudy (anonymous):

find an approximate solution of the logarithmic equation ln(x) +ln(x - 3) = 1

OpenStudy (loser66):

LHS ln a +ln b = ln(a*b) RHS 1 =e^0

OpenStudy (primeralph):